Show that the two curves (e', e², 1-e) and (1-theta , cos theta , sin theta ) intersect at the point (1, 1, 0). What is the angle between their tangents at that point?


1
Expert's answer
2021-09-26T15:47:09-0400

For the two curves to intersect at (1,1,0), then the value of each poianat of the curve must be the same and equal to (1,1,0)

Suppose "(e^t,e^{2t}, 1-e^t)=(1-\theta, \cos \theta, \sin \theta)\\
\implies t=0, \theta=0\\
\text{At } t=0\\
(e^t,e^{2t}, 1-e^t)=(1,1,0)\\
\text{At } \theta=0\\
(1-\theta, \cos \theta, \sin \theta)=(1,1,0)\\
\text{Hence, the two curves intersect at } (1,1,0).\\

\text{Let } r(t)=(e^t,e^{2t}, 1-e^t), ~~r(\theta)=(1-\theta, \cos \theta, \sin \theta)\\
r'(t)= (e^t, 2e^{2t}, -e^t)\\
r'(0)=(1,2,-1)\\
r'(\theta)=(-1, -\sin \theta, \cos \theta)\\
r'(0)=(-1,0,1)\\
\text{The angle between their tangent is }\\
\cos \alpha=\frac{(1,2,-1)\cdot(-1,0,1)}{(\sqrt{1^2+2^2+(-1)^2})\cdot \sqrt{(-1)^2+0^2+1^2}}\\
\cos \alpha=\frac{-2}{2\sqrt{3}}\\
\alpha=cos^{-1}(-\frac{1}{\sqrt{3}})\\
\alpha= 125.26°"